OpenStudy (anonymous):

ok im lost on trying this hw problem please explain how to do this (x-1)^2/5 = 4 im not sure how to factor the 2/5 power

OpenStudy (amistre64):

use logs: log(x-1^2/5) = log(4) 2/5 log(x-1) = log(4) log(x-1) = 5/2 log(4) log(x-1) = log(4^5/2) x-1 = 4^(5/2) x = 1 + 4^(5/2)

OpenStudy (amistre64):

or: 5root((x-1)^2) = 4 ^5 each side (x-1)^2 = 4^5 now square both sides x-1 = sqrt(4^5) x = 1+sqrt(4^5)

OpenStudy (amistre64):

x = 1 + sqrt(4*4*4*4*4) x = 1+5sqrt(4) x = 1+5(2) x = 11 by chance?

OpenStudy (amistre64):

i missed something in that...im sure of it :)

OpenStudy (anonymous):

this is suppose to be done by hand, i could easily throw it in the calculator and graph the equation for an answer. we have not learned anything about log yet awwwh that would work for me :) thx x is 32

OpenStudy (amistre64):

x = 33 ...maybe

OpenStudy (anonymous):

eeerr 33 i mean

OpenStudy (amistre64):

my 5sqrt(4) up there is a result of my own stupidity....just ignore the person behind the curtain :)

OpenStudy (anonymous):

lol thx yea i got it now i totally forgot how to do this

OpenStudy (amistre64):

what is the 5root(32^2) anyway...

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

4